OpenStudy (anonymous):

Select the equation that corresponds with the given graph This is a graph of a coordinate plane with a line. Two points of the line are zero comma five and one comma zero. y = -5x + 5 y equals negative 1 over 5 times x + 5 y equals 1 over 5 times x + 5 y = 5x + 5

OpenStudy (anonymous):

Are you single???

OpenStudy (anonymous):

no lol?

OpenStudy (johnweldon1993):

lol that was a slightly random response....regardless...back to the question lol you have 2 points (0 , 5) (1, 0) What you have already ...are the x and y intercepts y - intercept = 5 x - intercept = 1 so your equation will look something like y = mx + 5 We need to find the slope here use your 2 points (0 , 5) (1, 0) (x1,y1)(x2,y2) and this equation \[m = \frac{ y2 - y1 }{ x2 - x1 }\] to find the slope.. can you do this?

OpenStudy (anonymous):

I know haha but It would be the second or third choice, I believe

OpenStudy (johnweldon1993):

believe something else :P lol m = 0 - 5 ----- 1 - 0 m = -5 --- 1 m = -5 so your total equation would be y = 5x + 5 :)

OpenStudy (johnweldon1993):

edit *** y = -5x + 5

OpenStudy (anonymous):

What about Compare and Contrast. Below are two sets of points. Find the slopes between the two points and compare them. Choose the statement that is true about the values of the slopes. Set #1 Set #2 (1, 5) and (2, 9) (6, 2) and (3, 4) The slope of Set #1 is smaller than the slope of Set #2. The slope of Set #1 is larger than the slope of Set #2. The slope of Set #1 is the same as the slope of Set #2. None of the statements above describes slopes of sets shown.

OpenStudy (johnweldon1993):

Remember that equation I wrote up there? \[m = \frac{ y2 - y1 }{ x2- x1 }\] you have 2 points for each set set 1 set 2 (1, 5) and (2, 9) (6, 2) and (3, 4) (x1,y1) (x2,y2) (x1,y1) (x2,y2) do that equation for each set and what do you get for the slopes?

OpenStudy (anonymous):

Select the correct graph of -4x + 3y = 6. graph of line crossing the y axis at 2 and through point parentheses 3 comma 6 close parentheses graph of line crossing the y axis at 2 and through point parentheses 3 comma negative 2 close parentheses graph of line crossing the y axis at negative 2 and through point parentheses 3 comma 2 close parentheses graph of line crossing the y axis at negative 2 and through point parentheses 4 comma 5 close parentheses

OpenStudy (johnweldon1993):

So we would have the y-intercept....at 2 here and to check if the line goes through the point...you just plug in the point into the equation...and see if the equation is correct

OpenStudy (anonymous):

It would be option B?

OpenStudy (johnweldon1993):

Option A